# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ Abstract
Accurate orientation estimation is a crucial component of 3D molecular structure reconstruction, both in single-particle cryo-electron microscopy (cryo-EM) and in the increasingly popular field of cryo-electron tomography (cryo-ET). The dominant method, which involves searching for an orientation with maximum cross-correlation relative to given templates, falls short, particularly in low signal-to-noise environments. In this work, we propose a Bayesian framework to develop a more accurate and flexible orientation estimation approach, with the minimum mean square error (MMSE) estimator as a key example. This method effectively accommodates varying structural conformations and arbitrary rotational distributions. Through simulations, we demonstrate that our estimator consistently outperforms the cross-correlation-based method, especially in challenging conditions with low signal-to-noise ratios, and offer a theoretical framework to support these improvements. We further show that integrating our estimator into the iterative refinement in the 3D reconstruction pipeline markedly enhances overall accuracy, revealing substantial benefits across the algorithmic workflow. Finally, we show empirically that the proposed Bayesian approach enhances robustness against the “Einstein from Noise” phenomenon, reducing model bias and improving reconstruction reliability. These findings indicate that the proposed Bayesian framework could substantially advance cryo-EM and cryo-ET by enhancing the accuracy, robustness, and reliability of 3D molecular structure reconstruction, thereby facilitating deeper insights into complex biological systems.
@@ 1 Introduction
Determining the precise three-dimensional (3D) orientation of biological molecules from their noisy two-dimensional (2D) projection images is a fundamental challenge in cryo-electron microscopy (cryo-EM) [1, 30, 37, 7]. This process, known as orientation estimation, is crucial for various cryo-EM applications, including 3D reconstruction algorithms [44, 40], heterogeneity analysis [51, 55, 13, 63], and beyond [32]. Figure 1(a) illustrates the role of orientation estimation within the cryo-EM 3D reconstruction workflow.
@@
In cryo-electron tomography (cryo-ET), orientation estimation presents an additional challenge in the form of subtomogram averaging. Notably, cryo-ET suffers from higher noise levels compared to single-particle cryo-EM due to the complex and heterogeneous nature of cellular samples and the challenges of capturing data from multiple angles within thicker specimens. Subtomogram averaging offers an effective approach to enhance the signal-to-noise ratio (SNR), ultimately resulting in the reconstruction of high-resolution structures. This technique often involves extracting multiple similar subtomograms containing the target protein complex or macromolecule from a large cryo-electron tomogram reconstructed from all available tilts (typically from $-60^{\circ}$ to $+60^{\circ}$), followed by aligning and averaging them [61, 58]. Unlike traditional cryo-EM, this process typically aligns 3D structures directly, without the direct use of 2D projections (see Figure 1(b) for an illustration).
@@
Mathematically, the orientation estimation tasks in cryo-EM and cryo-ET are slightly different. In the process of cryo-EM, which involves 2D tomographic projections, the mathematical model can be formulated as:
$y=\Pi(g^{-1}\cdot V)+\varepsilon,$ (1.1)
@@
where $y$ is the observed 2D projection image, $V$ is the underlying 3D molecular structure, $\Pi$ is the tomographic projection operator, $g$ is the unknown 3D rotation operator of interest, $\varepsilon$ represents measurement noise, and $g^{-1}\cdot V(x)\equiv V(gx)$, representing a rotation acting on a volume $V$ with 3D coordinate $x$. Analogously, the mathematical model in cryo-ET subtomogram averaging, which aligns directly with the 3D structure, can be represented by:
$y=g^{-1}\cdot V+\varepsilon,$ (1.2)
@@
where $y$ is the observed 3D subtomogram, $V$ is the underlying 3D molecular structure, $g$ is the unknown 3D rotation operator of interest, and $\varepsilon$ represents measurement noise. Then, the goal of orientation estimation is to find the “best” 3D rotation $g$ based on the 2D projection image (in the cryo-EM case (1.1)) or the 3D subtomogram (in the cryo-ET case (1.2)) with respect to the 3D reference $V$. Namely, we aim to estimate the rotation $g$ given the sample $y$ and the 3D structure $V$.
While this work focuses on 3D rotation estimation, its methodology and results can be readily extended to various other transformations, including translations and permutations; we will discuss this further in Section 5.
1.1 The gap
@@
The common approach to estimating a rotation from an observation in the models above involves scanning through a pre-defined set of possible rotations and selecting the one that either maximizes the correlation or minimizes the distance to the given reference (2D projection image or 3D molecular structure, depending on the application). From the perspective of Bayesian statistics, this process corresponds to the maximum a posteriori (MAP) estimator, assuming a uniform prior over all the orientations (namely, all orientations are a priori equally likely). Therefore, we refer to this estimator as the MAP estimator in the sequel. However, Bayesian theory provides a much deeper and richer statistical framework that leads to improvement: replacing the MAP estimator with the Bayes estimator. The full potential of this estimator—that provides optimal accuracy according to a user-defined loss function and allows for integrating prior knowledge about rotational distributions—remained untapped so far.
The term rotational distribution broadly describes how rotations are distributed probabilistically over all possible orientations in 3D space. In cryo-electron microscopy, this concept becomes particularly relevant. During sample preparation, proteins are embedded in a thin layer of vitreous ice. Ideally, these proteins would adopt completely random orientations, resulting in what mathematicians refer to as a uniform distribution over the rotation group. In our model (1.1) or (1.2), this corresponds to the underlying prior distribution of the rotation $g$. This randomness is often assumed in modeling, as it simplifies analysis and reflects the desired experimental setup. However, in practice, proteins frequently display preferred specimen orientations due to interactions with the ice or other factors, leading to a distribution that is not uniform [53, 30, 29]. This means that some orientations are more likely than others, which has important implications for the methodologies employed in orientation estimation.
Currently, the vast majority of cutting-edge software for 3D structure reconstruction, e.g., RELION [44] and cryoSPARC [40], outputs the MAP orientation estimator. Although these tools are primarily designed for 3D structure reconstruction and treat rotation estimation as a nuisance problem, researchers may inadvertently use these orientation estimates for downstream tasks, potentially leading to suboptimal results. This work explores opportunities to improve orientation estimation, especially in low SNR conditions where the MAP estimator performs poorly.
1.2 The Bayesian framework
@@
The Bayesian framework has become a powerful and widely adopted tool in cryo-EM, now recognized as the leading method for recovering 3D molecular structures [44, 40, 55, 15]. It effectively addresses challenges like overfitting and parameter tuning while enhancing interpretability [43]. By explicitly modeling uncertainties, Bayesian methods enable more accurate and robust reconstructions of molecular structures, driving significant advances in both resolution and structural flexibility. While the majority of these methods focus on the task of structure reconstruction (e.g., see model (4.1)) and aim to achieve the MAP estimator for the volume structure [44, 40], the Bayesian framework offers broader possibilities for addressing other problems.
@@
In this work, we propose a versatile Bayesian framework specifically for orientation estimation, which has superior statistical properties and is highly flexible. This framework enables the use of any loss function over the rotational group tailored to users’ requirements and accommodates a broad range of prior distributions, beyond the uniform distribution. In particular, while the Bayesian framework can be adapted to different loss functions, this work focuses on the mean-squared error (MSE) loss, which is equivalent to the chordal distance between 3D rotations. The primary reason for using this loss function is that its corresponding Bayes estimator has a closed-form analytical solution, making it computationally efficient and easy to interpret. We denote this estimator as $\hat{g}_{\text{MMSE}}$, where MMSE stands for minimum mean square error. It is worth noting that for any given loss function and prior distribution, the Bayes estimator is optimal among all possible estimators. Moreover, the computational cost of calculating our $\hat{g}_{\text{MMSE}}$ estimator is at the same scale as that of the commonly used MAP orientation estimator (See Section 3 for more details).
@@
Figure 1 illustrates the critical role of orientation estimation in the cryo-EM and cryo-ET reconstruction processes, showcasing the superior performance of the MMSE orientation estimator, $\hat{g}_{\text{MMSE}}$, compared to the MAP orientation estimator, $\hat{g}_{\text{MAP}}$. Specifically, the curves show that the MMSE estimator consistently produces more accurate estimates than the MAP estimator, with the performance gap widening as the SNR decreases. Moreover, incorporating prior knowledge of rotational distribution into our MMSE estimator allows for even better performance (See Figure 2). This performance gap becomes even more pronounced when the estimators are incorporated into a reconstruction algorithm (see Section 4).
@@
The main takeaway of this paper is the recommendation to use the Bayesian MMSE orientation estimator to determine the orientation of each observation, in place of the commonly used MAP orientation estimator. The MMSE estimator demonstrates superior performance even under a uniform distribution, with the potential for further significant improvements when leveraging knowledge of underlying rotational distribution. Since the MMSE estimator for orientation determination is already integrated into many software packages, this adjustment should be easy to implement with minimal additional computational cost.
@@
Figure 1: Comparison between maximum cross-correlation ($\hat{g}_{\text{MAP}}$) and Bayesian ($\hat{g}_{\text{MMSE}}$) orientation estimators in cryo-electron microscopy (cryo-EM) and cryo-electron tomography (cryo-ET) applications. The figure illustrates the general workflow in cryo-EM and cryo-ET techniques, highlighting the role of orientation estimation in each technique. Panel (a) illustrates the model with 2D projections (cryo-EM model, (1.1)), while panel (b) shows the model without projections (cryo-ET model, (1.2)). (a) Cryo-EM involves imaging macromolecules embedded in a thin layer of vitreous ice using an electron beam in a transmission electron microscope (TEM). The process generates 2D projection images (micrographs) of particles in unknown 3D orientations. These 2D particles are then identified and extracted from the micrographs, forming the basis for subsequent steps of the macromolecule’s 3D structure reconstruction. (b) Cryo-ET involves imaging a sample from multiple known tilt angles (typically from $-60^{\circ}$ to $+60^{\circ}$) to create 2D projections, which are then combined computationally to reconstruct 3D subtomograms. In this context, a subtomogram refers to a small volume containing an individual 3D particle. The subtomograms are extracted by a particle picker algorithm for further analysis. (a+b): The rotation estimation problem involves determining the relative orientation of a noisy 2D particle (in cryo-EM) or a noisy 3D subtomogram (in cryo-ET) relative to a reference volume $V$. The reference volume structure used in both setups is identical and corresponds to the 80S ribosome [59]. Under high SNR conditions, both rotation estimators closely approximate the true relative rotation. However, as the SNR decreases, the estimation accuracy declines. Importantly, across all SNR levels, the geodesic angular distance between the MMSE orientation estimator and the true rotation consistently remains lower than that of the MAP orientation estimator. For (a), the estimation was conducted using a grid size of $L=3000$ samples of the rotation group $\mathsf{SO}(3)$, while for (b), a grid size of $L=300$ was used. Each point in the two curve plots represents the average error computed over 3000 trials.
1.3 Overview of results and contributions
Section 2 details the mathematical model and Bayesian approach, focusing on the MMSE estimator. The theoretical result examines the MMSE estimator’s behavior, explaining why it performs similarly to the MAP estimator in high-SNR conditions (Proposition 2.1), while consistently outperforming it in low-SNR conditions (Figure 1)—a frequent scenario in cryo-EM and cryo-ET applications [7].
Section 3 presents numerical simulations evaluating the performance of the MMSE estimator. These simulations explore the effects of various prior distributions on rotations and the influence of the sampling grid resolution, denoted as $L$, within the rotation group $\mathsf{SO}(3)$. Here, $\mathsf{SO}(3)$ refers to the special orthogonal group in three dimensions, representing the set of all possible 3D rotations. The parameter $L$ determines the granularity of the discretization applied to $\mathsf{SO}(3)$ in these simulations. We highlight that incorporating non-uniform prior distributions on rotations significantly improves estimation accuracy, emphasizing the importance of the prior information of rotational distribution and offering promising avenues for future integration within cryo-EM and cryo-ET frameworks (Figure 2). Moreover, empirical results show that in high-SNR conditions, the distance between the estimators and the true rotation scales proportionally to $L^{1/3}$ (Figure 3), whereas this effect diminishes in low-SNR conditions.
In Section 4, the MMSE estimator is connected to the orientation estimation step in 3D reconstruction algorithms (for models without projections), where soft-assignment methods are used. This connection demonstrates that incorporating the MMSE estimator into reconstruction algorithms is equivalent to the widely used expectation-maximization (EM) algorithm [46] (Proposition 4.1). Our experiments clearly show that both 2D and 3D reconstructions using the MMSE estimator significantly outperform those using the MAP estimator (Figures 4 and 5). Notably, the MMSE estimator demonstrates robustness against the “Einstein from Noise” phenomenon [19, 3], with empirical evidence showing reduced bias in 3D reconstruction (Figures 4 and 5). Section 5 summarizes the results and delineates future research directions.
1.4 Related Works
Estimating rotations in cryo-EM and cryo-ET involves two distinct but related tasks: 3D alignment and synchronization.
3D Alignment techniques.
3D alignment, which is a key focus of this work, aims to determine the relative orientation between two 3D volumes (in cryo-ET) or to estimate the orientation associated with a projection image (in cryo-EM). 3D alignment is often achieved through cross-correlation techniques, such as the MAP estimator. This task is commonly applied for downstream analysis or for comparing different conformational states.
Various alignment methods have been proposed in the literature. For instance, the Xmipp software package provides a 3D alignment algorithm based on spherical harmonics expansion and fast rotational matching [11]. The EMAN2 software package offers two 3D alignment methods: one that estimates rotation via an exhaustive Euler angle search with iterative angular refinement, and a more efficient tree-based approach that performs hierarchical 3D alignment with progressively finer downsampling in Fourier space [54]. Another approach is the projection-based volume alignment algorithm, which matches projections of the two volumes [60]. In [17], the authors propose leveraging common lines between projection images of the maps. Additionally,  [50] suggests that replacing the traditional Euclidean distance with the 1-Wasserstein distance creates a smoother optimization landscape and a larger basin of attraction.
Synchronization.
In contrast to 3D alignment, synchronization aims to recover a set of rotations from their corresponding rotated measurements without accessing the ground truth volumes. The standard approach follows a two-stage process: (1) estimating the relative rotation between each pair of observations; and (2) inferring global rotations from these pairwise ratios. The main techniques in this area rely on spectral decomposition [47], message passing [39], and convex optimization [5]. These methods have been applied to estimate projection orientations in cryo-EM, where the relative rotations between projections are determined using the common-lines property [56, 45, 49]. Recently, new deep learning-based methods have been developed for this task [36, 38, 24].
Rotationally invariant techniques.
It is worth noting that the unknown rotations in cryo-EM and cryo-ET are sometimes treated as nuisance parameters. This is because rotation estimation is often an intermediate step, with the primary goal being the estimation of the 3D structure itself. Consequently, some methods aim to estimate the 3D structure (or images) directly by leveraging features, such as polynomials, that are invariant to rotations (or translations). For example, in [62, 31], the bispectrum of images was used for rotationally invariant clustering. Similar techniques based on the method of moments have also been applied to 3D reconstruction; see, for example, [28, 8].

### v2

# Bayesian Perspective for Orientation Determination in Cryo-EM with Application to Structural Heterogeneity Analysis

@@ Abstract
Accurate orientation estimation is a crucial component of 3D molecular structure reconstruction, both in single-particle cryo-electron microscopy (cryo-EM) and in the increasingly popular field of cryo-electron tomography (cryo-ET). The dominant approach, which involves searching for the orientation that maximizes cross-correlation relative to given templates, is sub-optimal, particularly under low signal-to-noise conditions. In this work, we propose a Bayesian framework for more accurate and flexible orientation estimation, with the minimum mean square error (MMSE) estimator serving as a key example. Through simulations, we demonstrate that the MMSE estimator consistently outperforms the cross-correlation-based method, especially in challenging low signal-to-noise scenarios, and we provide a theoretical framework that supports these improvements.
When incorporated into iterative refinement algorithms in the 3D reconstruction pipeline, the MMSE estimator markedly improves reconstruction accuracy, reduces model bias, and enhances robustness to the “Einstein from Noise” artifact. Crucially, we demonstrate that orientation estimation accuracy has a decisive effect on downstream structural heterogeneity analysis. In particular, integrating the MMSE-based pose estimator into frameworks for continuous heterogeneity recovery yields accuracy improvements approaching those obtained with ground-truth poses, establishing MMSE-based pose estimation as a key enabler of high-fidelity conformational landscape reconstruction. These findings indicate that the proposed Bayesian framework could substantially advance cryo-EM and cryo-ET by enhancing the accuracy, robustness, and reliability of 3D molecular structure reconstruction, thereby facilitating deeper insights into complex biological systems.
@@ 1 Introduction
Determining the precise three-dimensional (3D) orientation of biological molecules from their noisy two-dimensional (2D) projection images is a fundamental challenge in cryo-electron microscopy (cryo-EM) bai2015cryo ; lyumkis2019challenges ; bendory2020single . This process, known as orientation estimation, is crucial for various cryo-EM applications, including 3D reconstruction algorithms scheres2012relion ; punjani2017cryosparc , heterogeneity analysis sorzano2019survey ; toader2023methods ; donnat2022deep ; zhong2021cryodrgn , and beyond maeots2022structural . For example, Figure 1(a) illustrates the role of orientation estimation within the cryo-EM 3D reconstruction workflow.
@@
In cryo-electron tomography (cryo-ET), orientation estimation presents an additional challenge in the form of subtomogram averaging. Notably, cryo-ET suffers from higher noise levels compared to single-particle cryo-EM due to the complex and heterogeneous nature of cellular samples and the challenges of capturing data from multiple angles within thicker specimens. Subtomogram averaging offers an effective approach to enhance the signal-to-noise ratio (SNR), ultimately resulting in the reconstruction of high-resolution structures. This technique often involves extracting multiple similar subtomograms containing the target protein complex or macromolecule from a large cryo-electron tomogram reconstructed from all available tilts (typically from $60^{\circ}$ to $60^{\circ}$), followed by aligning and averaging them zhang2019advances ; watson2024advances . Unlike traditional cryo-EM, this process typically aligns 3D structures directly, without the direct use of 2D projections (see Figure 1(b) for an illustration).
@@
Mathematically, the orientation estimation tasks in cryo-EM and cryo-ET are slightly different. In the process of single-particle cryo-EM, which involves 2D tomographic projections, the mathematical model can be formulated as:
$y=\Pi(g\circ V)+\varepsilon,$ (1.1)
@@
where $y:\mathbb{R}^{2}\to\mathbb{R}$ is the observed 2D projection image, $V:\mathbb{R}^{3}\to\mathbb{R}$ is the underlying 3D molecular structure, $\Pi$ is the tomographic projection operator, $g$ is the unknown 3D rotation operator of interest, $\varepsilon$ represents measurement noise, and $g\circ V(x)\equiv V(g^{-1}x)$, representing a rotation acting on a volume $V$ with 3D coordinate $x$. Analogously, the mathematical model in cryo-ET subtomogram averaging, which aligns directly with the 3D structure, can be represented by:
$y=g\circ V+\varepsilon,$ (1.2)
@@
where $y:\mathbb{R}^{3}\to\mathbb{R}$ is the observed 3D subtomogram, $V:\mathbb{R}^{3}\to\mathbb{R}$ is the underlying 3D molecular structure, $g$ is the unknown 3D rotation operator of interest, and $\varepsilon$ represents measurement noise. Then, the goal of orientation estimation is to find the “best” 3D rotation $g$ based on the 2D projection image (in the cryo-EM case (1.1)) or the 3D subtomogram (in the cryo-ET case (1.2)) with respect to the 3D reference $V$. Namely, we aim to estimate the rotation $g$ given the sample $y$ and the 3D structure $V$.
The actual mathematical models used in cryo-EM and cryo-ET are more intricate than those presented in (1.1) and (1.2), incorporating additional factors such as the contrast transfer function (CTF) and in-plane translations, as elaborated in Appendix A. Although the framework introduced in this work is capable of addressing the full cryo-EM model, we adopt the simplified models in (1.1) and (1.2) to more effectively communicate the primary insights.
The gap.
The common approach to estimating the rotation of an observation in the models above involves scanning through a pre-defined set of possible rotations and selecting the one that either maximizes the correlation or minimizes the distance to the given reference (2D projection image or 3D molecular structure, depending on the application). Typically, these metrics involve weighted correlations and distances, where the weights account for the noise characteristics scheres2012relion .
From an estimation theory perspective, this process corresponds to the maximum likelihood estimator (MLE), which does not incorporate prior information about the rotation distribution, i.e., the distribution of the 3D rotation $g$ appearing in models (1.1) or (1.2). A natural extension is the maximum a posteriori (MAP) estimator, which combines the data likelihood with a prior distribution on rotations. Notably, when a uniform prior is assumed, indicating that all orientations are equally likely a priori, the MAP estimator reduces to the MLE estimator steven1993fundamentals . A more rigorous treatment can be found in Section 2.3.
@@
However, Bayesian theory provides a much deeper and richer statistical framework that leads to improvement: replacing the MLE and MAP estimators with the Bayes estimator. The full potential of this estimator, that provides optimal accuracy according to a user-defined loss function and allows for integrating prior knowledge about rotation distribution, remained untapped so far.
The term rotation distribution refers to the probabilistic law governing how orientations are distributed over all possible rotations in 3D space. In cryo-EM, this concept is critical because molecules are first suspended in a thin aqueous solution and subsequently vitrified into amorphous ice, at which point their orientations become fixed but remain uncontrolled. Ideally, all orientations would be equally likely, corresponding to a uniform distribution. In practice, however, molecules often adopt preferred orientations, typically due to interactions with the air–water interface or other sample-preparation factors tan2017addressing ; lyumkis2019challenges ; li2021effect . As a result, certain orientations are disproportionately represented, introducing systematic biases that can degrade reconstruction quality and complicate orientation estimation. Accounting for these deviations from uniformity is therefore essential for both methodological development and the interpretation of cryo-EM data.
Currently, the most convenient way to obtain orientation estimates is through modules of reconstruction software such as RELION scheres2012relion and cryoSPARC punjani2017cryosparc , which output the MLE orientation estimator described earlier. These tools are primarily designed for 3D structure reconstruction and therefore treat orientation estimation as a nuisance problem; nevertheless, researchers frequently use the reported orientation estimates for downstream tasks such as heterogeneity analysis gilles2025cryo ; zhong2021cryodrgn and structural validation rosenthal2003optimal . However, relying on MLE-based estimates in these contexts can potentially lead to suboptimal results, especially under low-SNR conditions where MLE estimators are known to perform poorly. This work therefore explores opportunities to improve orientation estimation by exploiting Bayesian methods that incorporate prior knowledge of the rotation distribution.
The Bayesian framework.
@@
The Bayesian framework has become a powerful and widely adopted tool in cryo-EM, now recognized as the leading method for recovering 3D molecular structures scheres2012relion ; punjani2017cryosparc ; toader2023methods ; gilles2025cryo . It effectively addresses challenges like overfitting and parameter tuning while enhancing interpretability scheres2012bayesian . By explicitly modeling uncertainties, Bayesian methods enable more accurate and robust reconstructions of molecular structures, driving significant advances in both resolution and structural flexibility. While the majority of these methods focus on the task of structure reconstruction (e.g., see Section 4) and aim to achieve the MAP estimator for the volume structure scheres2012relion ; punjani2017cryosparc , the Bayesian framework offers broader possibilities for addressing other problems.
@@
This framework enables the use of any loss function over the rotational group tailored to users’ requirements and accommodates a broad range of prior distributions, beyond the uniform distribution. While the Bayesian framework can be adapted to different loss functions, this work focuses on the mean-squared error loss, which is equivalent to the chordal distance between 3D rotations. The primary reason for using this loss function is that its corresponding Bayes estimator has a closed-form analytical solution, making it computationally efficient and easy to interpret. We denote this estimator as $\hat{g}_{\text{MMSE}}$, where MMSE stands for minimum mean square error. It is worth noting that for any given loss function and prior distribution, the Bayes estimator is optimal among all possible estimators, in the sense that it minimizes the posterior expected loss. Moreover, the computational cost of calculating our $\hat{g}_{\text{MMSE}}$ estimator is at the same scale as that of the commonly used MLE and MAP orientation estimators (See Section 3 for more details).
Applications.
Orientation estimation is not only integral to 3D structure reconstruction but also underpins a range of downstream tasks. Below, we highlight several representative applications that demonstrate its broader methodological significance.
One prominent example is heterogeneity analysis, where the goal is to capture structural variability that may arise from differences in composition, discrete states, or continuous conformational changes. A substantial portion of state-of-the-art methods gilles2025cryo ; zhong2021cryodrgn ; levy2024end ; luo2023opus ; punjani2017cryosparc ; punjani20213d make the simplifying assumption that particle poses are already known and fixed, typically obtained from a consensus refinement procedure; these pose estimates then form the basis for subsequent modeling of structural variability. In this work, we consider continuous conformational heterogeneity and demonstrate in Section 5 that the quality of orientation estimation has a direct impact on the fidelity of the recovered conformational landscapes.
Another important application is 3D volume alignment. Unlike the setting of (1.2), where an observation is aligned to a known reference volume, this task involves aligning two noisy volumes of the same structure with unknown relative rotations. This scenario arises, for instance, in the computation of the Fourier Shell Correlation (FSC) curve, a standard tool for estimating the spectral signal-to-noise ratio (SSNR) and resolution. As shown in Appendix B, this problem can be expressed within the same statistical framework as (1.2), differing only by an additional noise term, which highlights its close methodological connection to orientation estimation.
A further noteworthy application concerns validation. A classic example is the method of tilt pairs introduced by Rosenthal and Henderson rosenthal2003optimal , in which a small subset of images (e.g., around ten) is collected at known relative tilt angles. While the absolute particle orientations remain unknown, the relative orientations between tilt pairs are specified by the experimental geometry. The degree of agreement between estimated and known relative rotations then provides a direct validation of orientation assignment accuracy.
These examples are by no means exhaustive, but they underscore that orientation estimation is not merely a technical nuisance in reconstruction workflows, but rather a key methodological component that enables diverse downstream analyses and validation strategies across cryo-EM and cryo-ET.
Overview of results and contributions.
In this work, we introduce a versatile Bayesian framework for orientation estimation with strong statistical guarantees and high flexibility. Section 2 formulates the problem and develops the Bayesian MMSE estimator. We show theoretically that the MMSE estimator coincides with the MLE estimator in high-SNR regimes (Proposition 2.2), while consistently outperforming it in low-SNR conditions (Figure 1), a typical scenario in cryo-EM and cryo-ET bendory2020single .
Section 3 evaluates the MMSE orientation estimator through simulations. We investigate the effect of different prior distributions and the discretization resolution $L$ of the rotation group $\mathsf{SO}(3)$. Results indicate that non-uniform priors substantially improve estimation accuracy, underscoring the value of incorporating prior knowledge (Figure 2). We further show that in high-SNR regimes the estimation error scales as $L^{1/3}$ (Figure 3), highlighting discretization as the dominant error source, whereas in low-SNR regimes noise dominates.
In Section 4, we connect the MMSE estimator to 2D image recovery and 3D structure reconstruction, showing that it naturally leads to the expectation-maximization (EM) framework for models without projections (Proposition 4.1). Empirical results confirm that the MMSE estimator consistently outperforms the MLE and MAP estimators in both tasks, providing higher accuracy and greater robustness against the “Einstein from Noise” artifact henderson2013avoiding ; balanov2024einstein (Figures 45).
Finally, Section 5 integrates MMSE-based pose estimation into structural heterogeneity analysis. Following the fixed-pose framework of RECOVAR, we show that MMSE pose estimator consistently improves the recovery of conformational variability compared to the MLE counterpart. Section 6 concludes with a discussion of further applications of the Bayesian approach to orientation estimation and future research directions.
Main takeaways.
@@
Figure 1 illustrates the critical role of orientation estimation in the cryo-EM and cryo-ET reconstruction processes, showcasing the superior performance of the MMSE orientation estimator, $\hat{g}_{\text{MMSE}}$, compared to the MLE orientation estimator, $\hat{g}_{\text{MLE}}$. Specifically, the curves show that the MMSE estimator consistently produces more accurate estimates than the MLE estimator, with the performance gap widening as the SNR decreases. Moreover, incorporating prior knowledge of rotation distribution into our MMSE estimator allows for even better performance (See Figure 2). This performance gap becomes even more pronounced when the orientation estimators are incorporated into a reconstruction algorithm (see Section 4).
Beyond reconstruction accuracy, we demonstrate that orientation estimation quality has a direct and substantial impact on downstream structural heterogeneity analysis. Understanding structural variability is essential for characterizing the dynamic behavior of macromolecular complexes. Methods such as RECOVAR gilles2025cryo , designed to recover structural heterogeneity, typically assume known particle poses (i.e., fixed-pose methods). In reality, these poses must be inferred, and MLE-derived estimates are commonly used in practice. We show that replacing MLE-derived poses with MMSE estimates in RECOVAR substantially improves the recovery of the latent conformational manifold, bringing results closer to those obtained using ground-truth poses. These findings establish MMSE orientation estimation not only as a tool that enhances reconstruction fidelity, but also as a facilitator that advances state-of-the-art continuous heterogeneity analysis.
@@
The main takeaway of this paper is the recommendation to adopt the Bayesian MMSE orientation estimator for determining the orientation of each observation, in place of the commonly used MLE estimator. The MMSE estimator demonstrates superior performance even under a uniform rotation distribution and offers further improvements when incorporating prior knowledge of the underlying rotation distribution. Importantly, while current software packages do not directly compute the MMSE estimator, they already calculate all the necessary components required for its implementation. Therefore, adopting the MMSE approach can be achieved with minimal additional computational cost. The detailed implementation and code are available at https://github.com/AmnonBa/bayesian-orientation-estimation.
@@
Figure 1: Comparison between maximum cross-correlation ($\hat{g}_{\text{MLE}}$) and Bayesian ($\hat{g}_{\text{MMSE}}$) orientation estimators in cryo-electron microscopy (cryo-EM) and cryo-electron tomography (cryo-ET) applications. The figure illustrates the general workflow in cryo-EM and cryo-ET techniques, highlighting the role of orientation estimation in each technique. Panel (a) illustrates the model with 2D projections (single-particle cryo-EM model, (1.1)), while panel (b) shows the model of the subtomogram averaging in cryo-ET, (1.2)). (a) Cryo-EM involves imaging macromolecules embedded in a thin layer of vitreous ice using an electron beam in a transmission electron microscope (TEM). The process generates 2D projection images (micrographs) of particles in unknown 3D orientations. These 2D particles are then identified and extracted from the micrographs, forming the basis for subsequent steps of the macromolecule’s 3D structure reconstruction. (b) Cryo-ET involves imaging a sample from multiple known tilt angles (typically from $60^{\circ}$ to $60^{\circ}$) to create 2D projections, which are then combined computationally to reconstruct 3D subtomograms. In this context, a subtomogram refers to a small volume containing an individual 3D particle. The subtomograms are extracted by a particle picker algorithm for further analysis. (a+b): The rotation estimation problem involves determining the relative orientation of a noisy 2D particle (in cryo-EM) or a noisy 3D subtomogram (in cryo-ET) relative to a reference volume $V$. The reference volume structure used in both setups is identical and corresponds to the 80S ribosome wong2014cryo . Under high SNR conditions, both rotation estimators closely approximate the true relative rotation. However, as the SNR decreases, the estimation accuracy deteriorates. Importantly, across all SNR levels, the geodesic angular distance between the MMSE orientation estimator and the true rotation consistently remains lower than that of the MLE orientation estimator. For (a), the estimation was conducted using a grid size of $L=3000$ samples of the rotation group $\mathsf{SO}(3)$, while for (b), a grid size of $L=300$ was used. Each point in the two curve plots represents the average error computed over 3000 trials.